OpenStudy (anonymous):

a divided by 6 times blank equals 5 times blank? what is the value of A???

OpenStudy (karatechopper):

ok well start by writing the equation out

OpenStudy (anonymous):

listen to kc

OpenStudy (anonymous):

Is it A/6 x ___ = 5 x ___ ? how is it possible to solve an equation with 2 unknowns?

OpenStudy (bahrom7893):

a/(6x)=5x It is possible jay

OpenStudy (bahrom7893):

u'll end up in terms of a

OpenStudy (bahrom7893):

multiply both sides by 6x, divide both sides by 30, that's your answer for x

OpenStudy (anonymous):

but i had thought they were asking for value of A... so value of A can refer to numerical or in terms of another unknown?

OpenStudy (bahrom7893):

well the question isn't worded properly.. is it: A/(6_) = 5_ or (A/6)_ = 5_

OpenStudy (bahrom7893):

in 2nd case a is just 6*5

OpenStudy (anonymous):

i wish i could give double medal to the same person...